This is a list, mostly for my self, of components used in the first filter cavity that are suboptimal, and should be replaced when the new components arive.
-1811 PD for REFL/PDH locking.
-PBS for power attenuation on 1811 PD (stolen from SHG sled). Causing power modulations when the fans are on.
-lots of BNC cables strung 2 or 3 together instead of a single long cable
Alignment is poor. I am still searching for flashes with the 775 light. The first pass beam can be seen with the NIR camera but nothing else.